Given the vertex and focus of a parabola, (4,2) and (4,5). Determine the following
a. General equation
b. Standard equation
c. Location of the opening
d. Location of the directrix (coordinate)
1
Expert's answer
2020-05-06T19:24:03-0400

Given the vertex and focus of a parabola, (4,2) and (4,5).

Since the x-coordinates of the vertex and focus are the same, so this is a regular vertical parabola, where the x part is squared. Since the vertex is below the focus, this is a right-side up parabola and p is positive. Since the vertex and focus are 5 – 2 = 3 units apart, then p = 3.

So the equation of parabola, we get

"(x-h)^2 = 4p(y-k)".

So, "(x-4)^2 = 12(y-2)".

Directrix equation is y = y-coordinate of vertex - p

So Directrix equation is "y = 2 - 3 = -1".


a) We found given parabola is

"(x-4)^2=12(y-2)"

So, General equation of parabola is: "x^2 - 8x - 12y + 40 = 0" .


b) Standard equation of parabola is

"(x-4)^2=12(y-2)".


c) Since the vertex is below the focus, this is a right-side up parabola.


d) Location of the Directrix: "y = -1."